# MAPS: A Multi-Agent Framework Based on Big Seven Personality and Socratic Guidance for Multimodal Scientific Problem Solving

### v1

@@ Abstract
Multimodal scientific problems (MSPs) involve complex issues that require the integration of multiple modalities, such as texts and diagrams, presenting a significant challenge in artificial intelligence. While progress has been made in addressing traditional scientific problems, MSPs still face two primary issues: the challenge of multi-modal comprehensive reasoning in scientific problem-solving and the lack of reflective and rethinking capabilities. To address these issues, we introduce a Multi-Agent framework based on the Big Seven Personality and Socratic guidance (MAPS) [^1] The source codes and experimental datasets are available online at https://github.com/exoskeletonzj/MAPS.. This framework employs seven distinct agents that leverage feedback mechanisms and the Socratic method to guide the resolution of MSPs. To tackle the first issue, we propose a progressive four-agent solving strategy, where each agent focuses on a specific stage of the problem-solving process. For the second issue, we introduce a Critic agent, inspired by Socratic questioning, which prompts critical thinking and stimulates autonomous learning. We conduct extensive experiments on the EMMA, Olympiad, and MathVista datasets, achieving promising results that outperform the current SOTA model by 15.84% across all tasks. Meanwhile, the additional analytical experiments also verify the model’s progress as well as generalization ability.
1 Introduction
@@
Multimodal scientific problems (MSPs) cover scientific scenarios that involve multiple modalities [5, 22, 16], such as text and vision. These problems typically span fields like mathematics, physics, and chemistry, requiring rigorous logical reasoning and solid domain expertise [3, 1, 13, 36, 41]. In the realm of artificial intelligence, effectively addressing these cross-modal, multi-domain challenges remains both important and difficult [12, 20]. Figure 1 illustrates a typical problem scenario that includes the context, the question statement, and an illustrative diagram. The diagram shows a lever divided into left and right sections: an unknown mass is hung on the left and a known mass on the right, each positioned at one-quarter of the total length from the respective ends, with a groove at the center. The accompanying description states that when the unknown mass is hung on the left, the system remains in static equilibrium. Solving this problem requires not only understanding the combined visual and textual information but also applying the lever balance principle by comparing the effects on both sides. One intuitive solution is to delegate the task-solving process to a single multimodal large language model (MLLM). Although the trending MLLMs have the basic abilities (e.g., diagram parsing and theorem retrieval), they are not well optimized to combine these skills in complex scenarios. Therefore, it motivates us to tackle this critical research question: how to leverage and elicit the off-the-shelf MLLMs to address the challenging MSPs?
To address the problem, this work proposes to establish a multimodal multi-agent framework for multimodal scientific problem-solving. Motivated by the Big Seven Personality theory (i.e., Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity) [2, 4, 31], the MLLMs can function as independent agents with specific skills. Within the framework, these agents are guided to perform the problem-solving in a deliberate, cooperative manner. As one of the initial attempts for MSPs, we offer a novel solution for handling multimodal scientific scenarios.
In solving MSPs, traditional approaches [35, 19, 15, 43, 6, 29] generally rely on a single MLLM to handle all tasks at once. Although these models possess basic abilities such as diagram parsing and theorem retrieval, they struggle to effectively integrate these skills in complex scenarios, resulting in what can be termed a multi-modal comprehensive reasoning challenge in scientific problem solving. This limitation arises because a single model cannot simultaneously manage diagram interpretation, context alignment, and professional knowledge supplementation, thus failing to mimic the step-by-step reasoning process of human cognition. In other words, while these models excel in specific tasks, they are unable to flexibly utilize various abilities to reason and solve complex multi-modal problems in a gradual, human-like manner. To address this, we propose a progressive four-agent solving strategy that sequentially employs a interpreter, an aligner, a scholar, and a solver, ultimately enhancing the accuracy and efficiency of solving multi-modal scientific problems.
The lack of reflective and rethinking capabilities is another issue in traditional MSPs solving methods. Single MLLMs typically perform one-time reasoning and lack the ability for repeated reflection, adjustment, and iteration, which are essential in human problem-solving. When solving complex problems, humans continuously refine and adjust their solutions based on initial reasoning results, gradually approaching the correct answer. In contrast, traditional models often remain stuck at the initial reasoning stage, unable to self-correct and optimize. To address this, we designed a Critic agent, inspired by Socratic questioning [10, 27], which has unique advantages in fostering critical thinking and stimulating autonomous learning. Through a heuristic feedback mechanism, the Critic provides multiple rounds of feedback and continuous correction throughout the four-step solving process, simulating the iterative reflection process of humans. This design effectively overcomes the limitations of single models in terms of reflective and iterative capabilities, improving the flexibility and accuracy of the solution process.
@@
Figure 2: The corresponding relation between the Big Seven Personality theory and the seven function-specific agents.
In summary, we propose a Multi-Agent framework for MSPs, based on the Big Seven Personality theory [2, 4, 31] and Socratic questioning (MAPS). This framework defines seven distinct roles—Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity—derived from the Big Seven theory, ensuring clear role delineation and mutual complementarity, as shown in Figure 2 and  Section 2.1. Together, these roles create a robust multi-agent architecture. Extensive experiments on the EMMA, Olympiad, and MathVista datasets show promising results across various reasoning benchmarks, confirming the effectiveness and superiority of our approach. Further analyses support our model’s advantages. The complete problem-solving process and data are provided in the supplementary materials. In summary, the contributions of this paper are threefold:
$\bullet$ We propose a multi-agent framework based on the Big Seven personality theory. The framework integrates different personality traits from the Big Seven model to create seven complementary agents. To the best of our knowledge, this represents the first attempt in the field.
$\bullet$ To tackle the challenges of multi-modal reasoning and iterative reflection, we propose a four-step strategy with a Critic agent inspired by Socratic questioning. This approach guides agents through structured stages, offering continuous feedback to refine reasoning.
$\bullet$ Extensive experiments on various scientific domain datasets demonstrate the effectiveness and superiority of the MAPS framework, offering 15.84% performance gains. Additional analytical experiments further validate the interpretability of both the optimization process and its results.
2 Methodology
In this section, we provide a detailed introduction to MAPS. Our design is based on Large Language Models (LLMs) as intelligent agents capable of performing various functions. We define the task of solving MSPs as follows: Given an input diagram $d_{i}$, a context $c_{i}$, and a question $q_{i}$, the expected output is the answer $a_{i}$, which can be either an option or a computed result. The formal definition of this process is:
$a_{i}=\mathcal{M}(d_{i},c_{i},q_{i}),$ (1)
where $\mathcal{M}(d_{i},c_{i},q_{i})$ is the response obtained when the input diagram $d_{i}$, context $c_{i}$, and question $q_{i}$ are processed together by the model to generate the answer $a_{i}$. The goal of this task is to provide accurate answers through the model.
Figure 3: The overall architecture of MAPS. It illustrates seven functional agents based on the Big Seven Personality theory. It first includes the Manager agent with predefined interaction logic and the UserProxy agent responsible for receiving user inputs. Subsequently, four specialized agents—Interpreter, Aligner, Scholar, and Solver—are introduced, each corresponding to a specific step in solving MSPs. Finally, the Critic agent is presented, providing feedback and corrections to ensure the results are more accurate and interpretable.
The model architecture for solving MSPs is shown in Figure 3. In Section 2.1, we introduce the seven distinct functional agents designed based on the Big Seven Personality theory. In Section 2.2, we present the predefined interaction logic in the Manager agent, along with the UserProxy agent responsible for receiving user input. In Section 2.3, we discuss the four specific agents—Interpreter, Aligner, Scholar, and Solver——each representing a step in solving MSPs. Finally, in Section 2.4, we introduce the Critic agent, which provides feedback and corrections for the specific solution steps to ensure the results are more accurate and interpretable.
2.1 Big Seven Personality
The Big Seven Personality theory is an extension of the traditional Big Five Personality [33, 8, 7], adding two other dimensions: Self-Esteem and Sensitivity. It divides human personality into seven main dimensions: Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity. Each reflects unique tendencies in emotions, cognition, and behavior, revealing the diversity among individuals in decision-making, communication, and problem-solving. This theory provides a systematic framework for understanding individual differences and offers valuable theoretical support and guidance for psychological research and practical applications.
Inspired by the Big Seven Personality theory, we designed a multi-agent system that strives to maintain both diversity and integrity in completing complex tasks. As shown in Figure 2, the seven personality traits correspond to seven agents, each performing different roles.
The Manager agent represents Conscientiousness, responsible for creating the experimental plan and schedule, ensuring each step is executed according to the predefined plan. The UserProxy agent represents Agreeableness, ensuring smooth information flow and coordinating the allocation of tasks within the experiment. The Interpreter agent represents Extraversion, responsible for interpreting diagrams into captions, thereby providing new ideas and information for MSPs to foster innovation. The Aligner agent represents Neuroticism, tasked with aligning the caption, context, and question to ensure the safe integration of these elements. The Scholar agent represents Openness, responsible for researching the professional knowledge required by MSPs and exploring various hypotheses. The Solver agent represents Self-Esteem, in charge of gathering all necessary information and resolving MSPs by selecting the most appropriate experimental approach. The Critic agent represents Sensitivity, providing feedback on the four solution steps performed by the Interpreter, Aligner, Scholar, and Solver agents, and offering evaluations along with adjustment strategies. The specific duties and role distribution of these agents are discussed in detail in the following sections.
2.2 Agents Interaction Logic
This section primarily introduces the interaction logic of the seven personality agents. Since solving MSPs requires multimodal semantic integration and multi-step reasoning within specialized domains, the Manager agent is responsible for formulating a detailed research plan and interaction logic to ensure the smooth execution of various system tasks. As shown in Figure 3, the Manager agent coordinates the operations of all other agents to ensure the task is completed successfully. Additionally, the UserProxy agent handles receiving user input, ensuring smooth communication between the user and the system, and facilitating effective interaction.
Algorithm 1 MAPS Procedure
@@ 2.2 Agents Interaction Logic
1:  Input: Diagram $d_{i}$, Context $c_{i}$, and Question $q_{i}$
2:  Output: Final answer $a_{i}$
3:  Manager: Formulate a detailed research plan
4:  UserProxy: Receive the input $x\in$ Dataset $D$
5:  Interpreter: Interpret the diagram into a caption $p_{i}=\mathcal{M}_{\text{int}}(d_{i})$
6:  Aligner: Align the caption with the question and context $l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i})$
7:  Scholar: Research for MSPs $s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i})$
8:  Solver: Solve based on the complete information $a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i})$
9:  Critic: Score and assess the four steps $r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i})$
10:  while True do
11:     if All Scores are 5 then Break
12:     else Choose the minimum score from $r_{i}$ and return to the step $\in\{\textit{Interpreter},\textit{Aligner},\textit{Scholar},\textit{Solver}\}$ to optimize the process.
13:     end if
14:  end while
15:  return  Final answer $a_{i}$
As illustrated in Algorithm 1, the entire workflow is orchestrated by the Manager agent, corresponding to the flowchart shown in Figure 3. Once the UserProxy agent receives the user input, the system proceeds through a four-step problem-solving process involving the Interpreter, Aligner, Scholar, and Solver agents. After completing these steps, the Critic agent, based on Socratic questioning, evaluates and scores the results of each step.
2.3 Four-Step MSPs Solving
This section will provide a detailed introduction to the functions of the Interpreter, Aligner, Scholar, and Solver agents.
Interpreter. To effectively integrate visual context into the textual context, we adopt a multi-modal fusion paradigm using the Interpreter agent. Initially, the Interpreter agent is employed to analyze diagrams meticulously, converting them into detailed and accurate captions. This process ensures that rich visual details are faithfully captured in the textual description, such as relative positions, dimensions, and other features. Subsequently, the generated caption is fused with the textual context and the question in later steps, achieving a cohesive integration of information that provides a richer and more consistent basis for problem-solving. The formula is shown as follows:
$p_{i}=\mathcal{M}_{\text{int}}(d_{i}),$ (2)
where $\mathcal{M}_{\text{int}}(\cdot)$ represents the Interpreter agent and $p_{i}$ denotes the caption.
Aligner. In textual form, the information conveyed by elements such as the context, questions, and options differs from the content presented in the diagram. These differences can lead to discrepancies in understanding the problem, which in turn may affect subsequent reasoning and judgment. Therefore, in order to gain a deeper understanding of the problem statement and accurately capture multimodal semantic information, precise alignment of these information granularities is crucial. As shown below, $l_{i}$ illustrates the result of the alignment between the diagram, context, and question:
$l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i}),$ (3)
where $\mathcal{M}_{\text{ali}}(\cdot)$ represents the aligner agent.
@@ 2.3 Four-Step MSPs Solving
Random Choice - 26.09 22.78 24.30 13.00 23.00 27.00 23.00 0.67 0.33 0.00 1.75 0.33 0.87 16.06
Human Expert - 56.09 55.74 55.90 75.00 64.50 86.00 64.50 48.00 34.67 30.36 54.17 12.33 37.80 52.73

### v2

# MAPS: Multi-Agent Personality Shaping for Collaborative Reasoning

@@ Abstract
Collaborative reasoning with multiple agents offers the potential for more robust and diverse problem-solving. However, existing approaches often suffer from homogeneous agent behaviors and lack of reflective and rethinking capabilities. We propose Multi-Agent Personality Shaping ((MAPS), a novel framework that enhances reasoning through agent diversity and internal critique. Inspired by the Big Five personality theory, MAPS assigns distinct personality traits to individual agents, shaping their reasoning styles and promoting heterogeneous collaboration. To enable deeper and more adaptive reasoning, MAPS introduces a Critic agent that reflects on intermediate outputs, revisits flawed steps, and guides iterative refinement. This integration of personality-driven agent design and structured collaboration improves both reasoning depth and flexibility. Empirical evaluations across three benchmarks demonstrate the strong performance of MAPS, with further analysis confirming its generalizability across different large language models and validating the benefits of multi-agent collaboration.
Codehttps://github.com/exoskeletonzj/MAPS
Introduction
Solving complex reasoning problems (bhattacharya2024large; li2024multimodal; he2024olympiadbench) often requires more than just accurate perception and factual recall—it demands nuanced interpretation, multi-step inference, and the ability to adapt when initial attempts fail (fu2024isobench; li2024survey). While large language models (LLMs) demonstrate promising capabilities in solving such problems, they frequently fall short in scenarios requiring sustained reasoning, internal verification, and flexible strategy revision (anand2024mm; alasadi2024multimodal; gao2024cantor; wang2024t; xu2025large). A key challenge lies in how to effectively combine diverse reasoning strategies and incorporate mechanisms for intermediate reflection. Prior work has explored both single-agent solutions and simple collaborative setups (e.g., paired discussion or voting) (kaesberg2025voting), yet these approaches often suffer from rigid behaviors and limited capacity for self-correction. This raises an important research question: how can we enhance the depth and adaptability of reasoning by enabling more flexible and reflective solution processes? Figure 1 illustrates a representative scenario that embodies these challenges, where successful problem solving requires interpreting multimodal input and applying domain-specific reasoning.
@@ Introduction
Figure 2: Comparison of reasoning paradigms. Single-agent and two-agent approaches offer limited adaptability. MAPS enables dynamic collaborative reasoning. Right: Built on GPT-4o, MAPS achieves the best performance across three benchmarks.
As illustrated in Figure 2, existing methods (wang2024examining; landau2024computational; hardiansyah2024analysis; zhang2024mm; caffagni2024revolution; qiu2024application) for complex reasoning problems often adopt single-agent solutions or simple two-agent collaborations. Although effective to some extent, these setups suffer from homogeneous agent behaviors—reasoning steps tend to repeat similar patterns, leading to redundancy and premature convergence. The lack of diversity limits exploration and reduces the chance of identifying alternative perspectives or correcting errors, especially in multi-turn settings where roles and strategies remain undifferentiated.
Another issue is the lack of reflective and rethinking capabilities in existing approaches. As shown in Figure 2, the interaction between two agents is often linear and lacks mechanisms for feedback or revision. Even with multiple turns, agents rarely revisit earlier reasoning or correct initial misconceptions. In contrast, human reasoning is inherently iterative: People reflect, reassess, and adjust their thinking over time. Without structured reflection, current methods risk premature convergence and fail to recover from early-stage errors.
To address these two issues, we propose MAPS (Multi-Agent Personality Shaping), a collaborative reasoning framework that enhances both diversity and adaptability in complex problem solving. Inspired by the Big Five personality theory (almagor1995big; benet1995big; simms2007big), MAPS shapes the reasoning behaviors of a set of role-specialized agents through distinct personality traits, promoting heterogeneous collaboration and mitigating behavioral homogeneity. To enable reflective thinking and iterative refinement, MAPS further introduces a Critic agent that revisits intermediate outputs, identifies flawed steps, and provides structured feedback. This integration of personality-guided agent design and internal critique supports deeper, more flexible reasoning aligned with human cognitive processes.
As shown in Figure 3, the interpreter (Openness) explores information from multiple angles, the aligner (Agreeableness) reconciles visual and textual cues, the scholar (Conscientiousness) ensures precision and factual grounding, the solver (Extraversion) drives goal-oriented conclusions, and the critic (Neuroticism) questions assumptions and detects flaws. The critic is further inspired by Socratic questioning (elder1998role; paul2019thinker), providing reflective feedback throughout the multi-stage reasoning process. Together, these roles enable a structured yet flexible collaboration that supports deeper and more reliable problem solving.
We conduct extensive experiments on three challenging benchmarks: MathVista (lu2023mathvista), OlympiadBench (he2024olympiadbench), and EMMA (hao2025can). These datasets cover a wide range of complex reasoning tasks. MAPS consistently outperforms baseline methods across all datasets, confirming its effectiveness in enhancing both reasoning depth and adaptability. We further evaluate MAPS under different base LLMs, and results consistently show its superiority across model backbones. Additional analyses examine the impact of the feedback mechanism and the overall efficiency of the framework.
Our main contributions are as follows:
$\bullet$ We propose MAPS, a multi-agent reasoning framework. To the best of our knowledge, this is the first work that incorporates personality shaping based on the Big Five theory into collaborative reasoning.
$\bullet$ MAPS addresses two key challenges in existing methods: behavioral homogeneity and lack of reflection. It assigns distinct personality traits to agents and introduces a Critic inspired by Socratic questioning.
$\bullet$ We conduct extensive experiments on three scientific reasoning benchmarks. MAPS achieves consistent performance gains (up to 15.84%) and generalizes well across different tasks and base models.
@@
Figure 3: The corresponding relation between the Big Five Personality theory and the five function-specific agents.
Methodology
Figure 4: The overall architecture of MAPS. The framework consists of five functional agents inspired by the Big Five personality theory. The core reasoning process is carried out by four specialized agents—Interpreter, Aligner, Scholar, and Solver—each responsible for a distinct stage in solving complex reasoning problems. Finally, the Critic agent provides reflective feedback and correction to enhance accuracy and interpretability.
This section introduces MAPS in four key components: Preliminaries, Agentic Interaction Logic, the Four-Step Reasoning process, and the Critic and Feedback mechanism.
Preliminaries
Task Definition.
We define the complex reasoning task as a mapping from structured multimodal inputs to a target answer space. Let $\mathcal{D}$, $\mathcal{C}$, and $\mathcal{Q}$ denote the input spaces of diagrams, contexts, and questions, respectively, and let $\mathcal{A}$ denote the answer space. Each instance is a triplet $(d_{i},c_{i},q_{i})\in\mathcal{D}\times\mathcal{C}\times\mathcal{Q}$, and the goal is to predict the corresponding answer $a_{i}\in\mathcal{A}$.
The task is thus modeled as a function:
$\mathcal{M}:\mathcal{D}\times\mathcal{C}\times\mathcal{Q}\rightarrow\mathcal{A},\quad a_{i}=\mathcal{M}(d_{i},c_{i},q_{i}),$ (1)
where $\mathcal{M}$ denotes the reasoning system responsible for processing visual-textual input and generating the output.
Agent-Based Modeling.
In MAPS, the reasoning process $\mathcal{M}$ is decomposed into a sequence of collaborative stages executed by a set of role-specialized agents $\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{K}\}$, where each agent $\mathcal{A}_{k}$ performs a specific function conditioned on both the original input and intermediate reasoning states.
Let $\mathcal{S}_{0}=(d_{i},c_{i},q_{i})$ be the initial input state. The overall reasoning unfolds as a staged transformation:
$\mathcal{S}_{k}=\mathcal{A}_{k}(\mathcal{S}_{k-1}),\quad\text{for }k=1,2,\dots,K,$ (2)
where $\mathcal{S}_{k}$ denotes the intermediate reasoning state after the $k$-th agent’s operation. The final answer is extracted as $a_{i}=\mathrm{Extract}(\mathcal{S}_{K})$. This formulation reflects the staged reasoning in MAPS, where specialized agents collaboratively refine the solution.
Agents Interaction Logic
As illustrated in Algorithm 1, this section introduces the interaction logic among the five personality-driven agents in MAPS. Complex problem reasoning requires multimodal semantic integration and structured inference over multiple steps. MAPS models this process as a functional composition of role-specialized reasoning agents, each shaped by a distinct personality trait.
Let $\mathbf{x}=(d_{i},c_{i},q_{i})$ be the input, and let $\mathbf{p}_{k}\in\mathbb{R}^{m}$ denote the personality embedding associated with agent $\mathcal{A}_{k}$. The entire collaborative reasoning process is represented as:
$a_{i}=\mathcal{F}(\mathbf{x};\mathbf{p}_{1},\dots,\mathbf{p}_{4})=\mathcal{A}_{4}\circ\mathcal{A}_{3}\circ\mathcal{A}_{2}\circ\mathcal{A}_{1}(\mathbf{x}),$ (3)
where each agent $\mathcal{A}_{k}(\cdot;\mathbf{p}_{k})$ executes its stage conditioned on its personality vector $\mathbf{p}_{k}$, contributing a unique reasoning perspective.
After the initial reasoning trajectory $\mathcal{T}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}\}$ is generated, the Critic agent performs a reflective evaluation by applying a feedback function:
$\mathbf{f}=\texttt{Reflect}(\mathcal{T}),\quad\mathbf{f}=\{f_{1},f_{2},f_{3},f_{4}\}\in\mathbb{R}^{4},$ (4)
where each score $f_{k}$ reflects the confidence of the $k$-th stage. If any $f_{k}<\tau$, the corresponding stage is revised. This feedback loop complements the forward reasoning path, allowing MAPS to emulate deliberative human cognition through structured collaboration and critique.
Proposition 1 (Monotonic Free-Energy Descent). For the free energy $F^{(t)}=\mathbb{E}_{q^{(t)}}[-\!\log p(\mathbf{x},a_{i}\mid\theta)]+\mathrm{KL}\bigl(q^{(t)}\|\;p\bigr)$, each Critic-triggered update satisfies
$F^{(t+1)}\;\leq\;F^{(t)}.$ (5)
Thus the MAPS iteration produces a non-increasing free-energy sequence that converges to a stationary point. The full proof is provided in Appendix A.1.
Algorithm 1 MAPS Collaborative Reasoning with Reflective Feedback
@@
1:Input: Diagram $d_{i}$, Context $c_{i}$, Question $q_{i}$, Personality vectors $\{\mathbf{p}_{1},\dots,\mathbf{p}_{4}\}$
2:Initialize: $\mathcal{S}_{0}=(d_{i},c_{i},q_{i}),\quad t\leftarrow 0$
3:repeat
4:  Interpreter:  $p_{i}=\mathcal{A}_{1}(\mathcal{S}_{0};\mathbf{p}_{1})$
5:  Aligner:  $l_{i}=\mathcal{A}_{2}(p_{i},c_{i},q_{i};\mathbf{p}_{2})$
6:  Scholar:  $s_{i}=\mathcal{A}_{3}(l_{i},p_{i},c_{i},q_{i};\mathbf{p}_{3})$
7:  Solver:  $a_{i}=\mathcal{A}_{4}(s_{i},l_{i},p_{i};\mathbf{p}_{4})$
[…]
9:  Critic:  $\mathbf{f}^{(t)}=\texttt{Reflect}(\mathcal{T}^{(t)}),\quad f_{k}\in[0,1]$
10:  if all $f_{k}\geq\tau$ then
11:   break
12:  else
13:   Identify stage $k^{*}=\arg\min_{k}f_{k}$
14:   Rerun agent $\mathcal{A}_{k^{*}}$ with updated input
15:   $t\leftarrow t+1$
16:  end if
17:until convergence
18:return Final answer $a_{i}$
Four-Step Reasoning
Given input $\mathbf{x}=(d_{i},c_{i},q_{i})$, MAPS conducts a structured reasoning process across four personality-driven agents: Interpreter, Aligner, Scholar, and Solver. Each agent $\mathcal{A}_{k}$ is parameterized by a personality embedding $\mathbf{p}_{k}\in\mathbb{R}^{m}$, shaping its reasoning style and focus. The entire inference pipeline can be expressed as a function composition:
$a_{i}=\mathcal{A}_{4}\circ\mathcal{A}_{3}\circ\mathcal{A}_{2}\circ\mathcal{A}_{1}(\mathbf{x};\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3},\mathbf{p}_{4}),$ (6)
where each intermediate state $\mathcal{S}_{k}$ is defined recursively as $\mathcal{S}_{k}=\mathcal{A}_{k}(\mathcal{S}_{k-1};\mathbf{p}_{k})$, with $\mathcal{S}_{0}=\mathbf{x}$.
Interpreter. The Interpreter agent aims to extract structured visual semantics from diagram $d_{i}$, translating them into a caption representation $p_{i}$ that can be consumed by downstream language agents. Let $\phi_{\text{vis}}(d_{i})$ be a visual encoder, and $\psi_{\text{lang}}(\cdot)$ a caption generator. Then the agent performs:
$p_{i}=\mathcal{A}_{1}(d_{i};\mathbf{p}_{1})=\psi_{\text{lang}}\big(\phi_{\text{vis}}(d_{i})+W_{1}\mathbf{p}_{1}\big),$ (7)
where $W_{1}\in\mathbb{R}^{d\times m}$ projects the personality embedding into the encoder space, modulating its attention to attributes like spatiality, color, or topology.
Aligner. To resolve semantic mismatches across modalities, the Aligner fuses the interpreted diagram caption $p_{i}$ with the textual context $c_{i}$ and question $q_{i}$. The process outputs an alignment representation $l_{i}$, optimized to preserve shared semantics and suppress modality conflict. Formally,
$l_{i}=\mathcal{A}_{2}(p_{i},c_{i},q_{i};\mathbf{p}_{2})=\mathrm{CrossFuse}\big(p_{i},c_{i},q_{i};\mathbf{p}_{2}\big),$ (8)
where $\mathrm{CrossFuse}(\cdot)$ denotes a multi-head attention-based fusion operator, adaptively weighted by $\mathbf{p}_{2}$ to emphasize visual or linguistic cues based on agent bias.
Scholar. While $l_{i}$ captures semantic consistency, complex reasoning often requires external knowledge supplementation. The Scholar agent retrieves and integrates domain-specific knowledge $\mathcal{K}(l_{i})$, such as physics principles or mathematical theorems. We define:
$s_{i}=\mathcal{A}_{3}(l_{i},p_{i},c_{i},q_{i};\mathbf{p}_{3})=\mathrm{KnowAug}\big(l_{i},\mathcal{K}(l_{i});\mathbf{p}_{3}\big),$ (9)
where $\mathrm{KnowAug}(\cdot)$ augments contextual embeddings with retrieved tuples $\mathcal{K}(l_{i})$ from a structured knowledge memory, and $\mathbf{p}_{3}$ biases the agent toward formal rigor or heuristic reasoning.
Solver. The Solver agent aggregates all upstream outputs and executes logical composition to generate the final answer $a_{i}$. Let $\mathcal{H}_{i}=\{p_{i},l_{i},s_{i}\}$ be the hybrid reasoning state. The solver computes:
$a_{i}=\mathcal{A}_{4}(\mathcal{H}_{i};\mathbf{p}_{4})=\mathrm{Deduct}\big(p_{i},l_{i},s_{i};\mathbf{p}_{4}\big),$ (10)
where $\mathrm{Deduct}(\cdot)$ is a constrained generation module that synthesizes the inputs under logical, numerical, or symbolic rules. The final prediction $a_{i}\in\mathcal{A}$ may be a selected option or a free-form answer.
Together, these four stages construct a trajectory $\mathcal{T}=\{p_{i},l_{i},s_{i},a_{i}\}$, on which the Critic agent operates for reflection and feedback.
Proposition 2 (Collaborative Information Bottleneck). Let $\mathcal{S}_{k}$ be the intermediate output of the $k$-th agent, given input $\mathbf{x}=(d_{i},c_{i},q_{i})$ and target answer $a_{i}$. Then the MAPS reasoning process optimizes
$\min\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad I(\mathcal{S}_{k};a_{i})\geq\varepsilon,$ (11)
where $I(\cdot;\cdot)$ denotes mutual information. The Critic monitors constraint violations and reactivates stages with insufficient task-relevant information.
A derivation and discussion are provided in Appendix A.2.
@@
Random Choice - 26.09 22.78 24.30 13.00 23.00 27.00 21.00 0.67 0.33 0.00 1.75 0.33 0.87 16.06
Human Expert - 56.09 55.74 55.90 75.00 64.50 86.00 75.17 48.00 34.67 30.36 54.17 12.33 37.80 52.73